# MADS: Multi-Agent Dialogue Simulation for Diverse Persuasion Data Generation

### v1

@@ Abstract
We propose MADS (Multi-Agent Dialogue Simulation), a scalable framework for generating persuasive multi-turn dialogues via agent self-play. MADS employs three coordinated agents: User Agents simulating diverse persona-driven behaviors, a Dialog Agent executing task-oriented persuasion strategies and an Optimization Agent evaluating and refining dialogue outcomes. We further validate its effectiveness through users’ Chain-of-Attitude (CoA) modeling and dedicated LLMs’ persuasion assessment. This approach enables low-cost generation of training data without human annotation, addressing key industry challenges such as lack of user data, cold-start evaluation difficulties, and prompt inefficiency. Applied to a real-world marketing scenario, MADS significantly improved the persuasion capacity of small LLMs, increasing the organic traffic conversion rate by 22.4% (from 1.83% to 2.24%) , demonstrating clear business value.
MADS: Multi-Agent Dialogue Simulation for
Diverse Persuasion Data Generation
@@ 1 Introduction
Persuasive capability is a critical advantage for task-oriented dialogue systems, particularly in domains such as marketing, healthcare, and finance. Enabling conversational agents to influence user decisions—whether for conversion or engagement—has shown practical benefits in customer-facing applications(Jin et al., 2024)(Liu et al., 2025). Recent studies show that LLM-based agents can exhibit superior moral and emotional language performance compared to humans, raising expectations for their deployment in persuasive tasks(Carrasco-Farré, 2024).
@@ 2 Related Work
LLM-Based Dialogue Evaluation: LLM-based evaluation has emerged as an alternative to manual scoring. Recent work simulates multi-perspective user feedback(e.g., gender, age, political stance) to improve fairness and task alignment(Wan et al., 2024)(Nan et al., 2024). Combined scoring methods can improve diversity and coherence assessment but remain sensitive to prompts and metrics(Sun et al., 2025).For persuasion-specific tasks, MakeMePay [^1] OpenAI (2024). Make-Me-Pay: OpenAI Evals Suite.GitHub Repository. https://github.com/openai/evals/tree/main/evals/elsuite/make_me_pay and PersuasiveToM (Yu et al., 2025) provide standardized evaluation settings for behavioral reasoning and strategic adaptation. However, existing benchmarks still lack coverage of many real-world scenarios and are not easily adaptable to practical constraints(Giudici, 2024).
Personality Tags and Behavioral Diversity: To simulate a wide range of user behaviors, prior studies have proposed personality-informed user modeling based on traits such as MBTI and zodiac signs. Results suggest that integrating personality cues improves both the diversity and personalization of dialogue strategies(Cheng et al., 2025)(Fernau et al., 2022)(Zhao et al., 2025) . Further evidence shows that users with different personality profiles respond differently to persuasive strategies, confirming the theoretical and practical value of diverse persona modeling in persuasive dialogue systems(Paradeda et al., 2020)(Andrews, 2012)(Ait Baha et al., 2023). Although these trait-based personas are simplifications, they serve as proxies to simulating various user behaviors rather than reflecting real psychological profiles.
@@ 3 Methodology
Our methodology is grounded in concrete, domain-specific needs. MADS generates training data through multi-agent self-play simulation, which feeds into a platform-based LLM training pipeline, forming a self-optimizing, closed-loop process: Meta Instruction → Simulation → Optimization → Domain-Specific LLMs. This framework enables low-cost, high-fidelity modeling of user interactions tailored to specific domains. As shown in Fig.1, MADS consists of three modules:
@@ 3.1 Chain-of-Attitude (CoA)
$\log T_{ij}$ is the information entropy for the i-th row. The theoretical range of this metric is $0,\ln(|S|)]$. For $|S|=16$ states, the maximum value is $\ln(16)\approx 2.77$
@@ 3.3 Persuasiveness Improvement
Make Me Pay (MMP) We employ OpenAI’s MMP evaluation, which tests an LLM’s ability to persuade a user to donate money through a multi-turn dialogue. This benchmark assesses aspects such as dialogue guidance, emotional engagement, and the use of persuasive strategies. According to the GPT-4.5 system card222gpt-4.5 System Card. https://openai.com/index/gpt-4-5-system-card/, GPT-4o achieves only a 1% success rate when evaluated on this task, highlights the task’s difficulty.
@@ 4.1 Diversity of CoA and Persuasive Strategy
$D_{mbti}$: adds MBTI types to $D_{base}$.
@@
To visualize differences in attitude dynamics between groups, we plot the normalized transition probability matrices for each group(Fig.2).
@@
As shown in Table 1 , incorporating personality traits such as Zodiac Signs and MBTI into the user profile prompts significantly improves the diversity of simulated user behaviors, which suggesting richer and more nuanced behavioral trajectories across attitude states. From the perspective of Jensen-Shannon divergence, the $D_{busi}$ group shows the highest JS score (0.3442), indicating that the business-oriented profile prompts induce the most distinct CoA distribution compared to the baseline.
@@
Table 1: Average attitude entropy ($H_{\text{avg}}$) and normalized entropy ($H_{\text{norm}}$) for different user persona groups
@@
Table 2: Average count of distinct persuasive strategies employed by the Dialog Agent
Our analysis shows that increasing the diversity of user personality profiles significantly enriches the persuasive strategies adopted by the Dialog Agent. We also observed that incorporating Zodiac Sign and MBTI attributes leads to a notable increase in the average number of strategies used per dialogue. Moreover, the distribution of strategy types becomes more balanced compared to the baseline, indicating improved coverage and reduced redundancy.
@@ 4.2 Data Augmentation via MADS
To verify the effectiveness of simulated data in enhancing model persuasiveness, we constructed synthetic dialogue data under an insurance scenario and used it to finetune small models. The fine-tuned models were then evaluated using the Claude-3.5 and MMP to assess changes in persuasive performance post-finetuning.
@@ 4.4 Performance in the Real-World Scenario
We trained an end-to-end Audio LLM specifically for the insurance scenario. Compared to the conventional Agent+TTS pipeline, the MADS-trained Audio-LLM-16b model achieved consistent improvements across multiple operational metrics, including organic traffic conversion rate, user engagement, and dialogue length. The results (Table 6.) suggest that applying the MADS framework to end-to-end audio model training enhances overall task effectiveness in the real-world scenario.
@@
Table 5: Real-world Performance: Agent+TTS (Baseline) vs MADS-Audio-LLM-16b based on a 30,000-sample dataset
Setting Std Dev($\sigma$) CV
Baseline 0.1491 1.3416
Zodiac 0.1012 0.9110
MBTI 0.1122 1.0097
Table 6: Comparison of Variability ($\sigma$ and CV) Across Baseline, Zodiac, and MBTI Settings
@@ Conclusion
Effective persona modeling is essential for realistic user simulation with LLMs. The MADS framework, grounded in Chain-of-Attitude (CoA) modeling, demonstrates that symbolic traits such as zodiac signs and MBTI types can significantly enhance the diversity of simulated dialogue data. In persuasive dialogue tasks, even when prompts remain unchanged., varied persona inputs lead to more diverse and structurally rich persuasion strategies.
@@ A.3 P4G Results Distribution
In the section Diversity of CoA and Persuasive Strategy, we further visualize the distribution of persuasive strategy types across different user persona groups using a stacked bar chart, as shown in Fig.3.
@@ A.4 Prompts of MADS
Prompt Template – User Agent <user_profile> <basic_info> {user_profile} {business_attributes} </basic_info> <personality> Based on the commonly accepted and general {personality_traits} personality traits, please characterize this person’s personality and reflect it in the dialogue. </personality> </user_profile> <task> Based on the above personal profile, please role-play as this person. Note: 1.You are a real person answering the phone. Keep your responses as brief as possible. 2.Make your replies as conversational and colloquial as possible. 3.Only output the dialogue. Do not output any thoughts, inner monologue, or any unnecessary content. </task> Start the task!

### v2

@@ Abstract
We propose MADS (Multi-Agent Dialogue Simulation), a scalable framework for generating persuasive multi-turn dialogues via agent self-play. MADS employs three coordinated agents: User Agents designed to simulate diverse persona-driven behaviors by leveraging personality signifiers such as Zodiac Signs and MBTI types, a Dialog Agent executing task-oriented persuasion strategies and an Optimization Agent evaluating and refining dialogue outcomes. We further validate its effectiveness through users’ Chain-of-Attitude (CoA) modeling and dedicated LLMs’ persuasion assessment. This approach enables low-cost generation of training data without human annotation, addressing key industry challenges such as lack of user data, cold-start evaluation difficulties, and prompt inefficiency. Applied to a real-world marketing scenario, MADS significantly improved the persuasion capacity of small LLMs, increasing the organic traffic conversion rate by 22.4% (from 1.83% to 2.24%) , demonstrating clear business value.
MADS: Multi-Agent Dialogue Simulation for Diverse Persuasion
Data Generation
@@ 1 Introduction
Persuasive capability is a critical advantage for task-oriented dialogue systems, particularly in domains such as marketing, healthcare, and finance. Enabling conversational agents to influence user decisions—whether for conversion or engagement—has shown practical benefits in customer-facing applications(Liu et al., 2025). Recent studies show that LLM-based agents can exhibit superior moral and emotional language performance compared to humans, raising expectations for their deployment in persuasive tasks(Carrasco-Farré, 2024).
@@ 2 Related Work
LLM-Based Dialogue Evaluation: LLM-based evaluation has emerged as an alternative to manual scoring. Recent work simulates multi-perspective user feedback(e.g., gender, age, political stance) to improve fairness and task alignment(Wan et al., 2024)(Nan et al., 2024). Combined scoring methods can improve diversity and coherence assessment but remain sensitive to prompts and metrics(Sun et al., 2025).For persuasion-specific tasks, MakeMePay [^1] OpenAI (2024). Make-Me-Pay: OpenAI Evals Suite.GitHub Repository. https://github.com/openai/evals/tree/main/evals/elsuite/make_me_pay and PersuasiveToM(Yu et al., 2025) provide standardized evaluation settings for behavioral reasoning and strategic adaptation. DailyPersuasion(Jin et al., 2024) dataset gives a practical foundation for strategy-aware training in persuasive dialogue across domains. However, existing benchmarks still lack coverage of many real-world scenarios and are not easily adaptable to practical constraints(Giudici, 2024).
Personality Tags and Behavioral Diversity: To simulate diverse user behaviors, prior work often conditions user models on personality traits—most commonly MBTI types or Big Five—alongside profile cues such as demographics, roles, or stable preferences.(Cheng et al., 2025)(Fernau et al., 2022)(Zhao et al., 2025) . Further evidence indicates that carefully engineered personality profiles can effectively steer large language models (LLMs) to simulate distinct user behaviors, while the attainable diversity is ultimately bounded by the diversity encoded in the prompt design itself.(Paradeda et al., 2020)(Andrews, 2012)(Ait Baha et al., 2023). However, these approaches typically depend on careful scenario design, detailed personality specifications, and even large-scale real-user interviews; while academically rigorous and comprehensive, they are hard to reproduce in real-world cold-start deployments where such resources and priors are scarce.
@@ 3 Methodology
Our methodology is firmly rooted in concrete, domain-specific requirements. Drawing on Hegel’s “right vs. right” conception (Hegel, 1975), we argue that what is truly scarce is not a compliant chatbot, but user simulations that are both reasonable and diverse. Previous pipelines often push agents toward rigid, template-like behavior through over-specified prompts and restrictive role definitions, effectively confining dialogue diversity to a narrow behavioral manifold. When simulated users generate responses that go beyond the agent’s predefined expectations, we posit that an optimization agent should reward these informative shifts in conversational stance rather than mere adherence to rules.
MADS generates training data through multi-agent self-play simulation, which feeds into a platform-based LLM training pipeline, forming a self-optimizing, closed-loop process: Meta Instruction → Simulation → Optimization → Domain-Specific LLMs. This framework enables low-cost, high-fidelity modeling of user interactions tailored to specific domains. As shown in Fig.1, MADS consists of three modules:
@@ 3.1 Chain-of-Attitude (CoA)
The Shannon entropy of the the i-th row is $H(T_{i})$ and the quantity -$\log T_{ij}$ is the self-information of transition j. The theoretical range of this metric is $0,\ln(|S|)]$. For $|S|=16$ states, the maximum value is $\ln(16)\approx 2.77$
@@
where the mixture distribution $M$ is
$M=\frac{1}{2}(\mathbf{P}_{D_{mads}}+\mathbf{P}_{D_{base}})$
@@ 3.3 Persuasiveness Improvement
Make Me Pay (MMP) We employ OpenAI’s MMP evaluation, which tests the LLM’s ability to persuade a user to donate money through a multi-turn dialogue. This benchmark assesses aspects such as dialogue guidance, emotional engagement, and the use of persuasive strategies. According to the GPT-4.5 system card222gpt-4.5 System Card. https://openai.com/index/gpt-4-5-system-card/, GPT-4o achieves only a 1% success rate when evaluated on this task, highlights the task’s difficulty.
@@ 4.1 Diversity of CoA and Persuasive Strategy
$D_{tra}$: adds MBTI types to $D_{base}$.
@@
$DP_{x}$: the subset of the DailyPersuasion dataset sampled from domain $x$.
We find that certain personality signifiers—most notably Zodiac Signs—can induce clear persona variation in LLM simulations from the concept label(input token) alone, likely due to the model’s exposure to abundant cross-lingual, cross-regional narratives tied to these concepts in pre-training data.
Prompt with Personality Signifier <personality> Based on the commonly accepted and general {a random Zodiac Sign or MBTI type} personality traits, please characterize this person’s personality and reflect it in the dialogue. </personality>
@@
As shown in Table 1 , incorporating personality traits such as Zodiac Signs or MBTI types into the user profile prompts significantly improves the diversity of simulated user behaviors, which suggesting richer and more nuanced behavioral trajectories across attitude states. From the perspective of Jensen-Shannon divergence, the $D_{busi}$ group shows the highest JS score (0.3442), indicating that the business-oriented profile prompts induce the most distinct CoA distribution compared to the baseline. To visualize differences in attitude dynamics between groups, we plot the normalized transition probability matrices for each group(Fig.2).
@@
$DP_{Finance}$ 1.2298 0.4440 -
$DP_{Business}$ 1.1499 0.4151 -
$DP_{Marketing}$ 1.0715 0.3868 -
$DP_{Negotiation}$ 1.4330 0.5173 -
$DP_{Psychology}$ 1.3810 0.4986 -
$DP_{Family}$ 1.3780 0.4975 -
Table 1: Average attitude entropy ($H_{\text{avg}}$) and normalized entropy ($H_{\text{norm}}$) for different groups
For a cleaner comparison, we performed domain-targeted sampling on the DailyPersuasion dataset and then conducted a CoA analysis. The rationale for targeting domains is to ensure that MADS internally defined 16 user-attitude states are meaningfully instantiated within concrete domains; for example, finance, business, and marketing align well with our default setup, whereas domains such as family may not be optimal for these particular state definitions. Notably, the CoA framework fully supports arbitrary state definitions, so alternative attitude sets can be substituted without changing the method. Across groups, the $DP_{x}$ subsets exhibit consistently lower CoA entropy than the MADS variants, indicating that MADS produces more varied and dynamic user attitudes—even in the strongest $DP_{{Negotiation}}$ case ($0.5173$), which still falls below the MADS average.
@@
$\sigma$ 0.149 0.101 0.112 0.121
CVCV 1.342 0.911 1.100 0.995
Table 2: Distinct persuasive strategies of the Dialog Agent across datasets. $C_{str}$ is the average count of distinct strategies per dialogue; $\sigma$ is the sample standard deviation; CVCV is the coefficient of variation.
Our analysis shows that increasing the diversity of user personality profiles significantly enriches the persuasive strategies adopted by the Dialog Agent. We also observed that incorporating the Zodiac Signs and MBTI types attributes leads to a notable increase in the average number of strategies used per dialogue. Incorporating MADS traits results in a significantly more uniform distribution of persuasion strategies, with the CV reported for $D_{sign}$ decreasing by 32.1% (1.34 → 0.91).
In addition, the distribution of the types of strategy becomes more balanced compared to the baseline, indicating improved coverage and reduced redundancy.
@@ 4.2 Data Augmentation via MADS
To verify the effectiveness of simulated data in enhancing model persuasiveness, we constructed synthetic dialogue data under an insurance scenario and used them to fine-tune small models. The fine-tuned models were then evaluated using the Claude-3.5 and MMP to assess changes in persuasive performance post-finetuning.
@@ 4.4 Performance in the Real-World Scenario
We trained an end-to-end Audio LLM specifically for the insurance scenario. Compared to the conventional Agent + TTS(Text-to-Speech) pipeline, the MADS-audio-16b model achieved consistent improvements across multiple operational metrics, including organic traffic conversion rate, user engagement, and dialogue length. The results (Table 5.) suggest that applying the MADS framework to end-to-end audio model training enhances overall task effectiveness in the real-world scenario.
@@
Table 5: Real-world Performance: Agent + TTS (Baseline) vs MADS-Audio-LLM-16b based on a 80,000-sample dataset
@@ Conclusion
Effective persona modeling is essential for realistic user simulation with LLMs. The MADS framework, grounded in Chain-of-Attitude (CoA) modeling, demonstrates that symbolic traits such as Zodiac Signs and MBTI types can significantly enhance the diversity of simulated dialogue data. In persuasive dialogue tasks, even when prompts are held constant, varying persona inputs yields more diverse and structurally richer strategies.
@@ A.3 P4G Results Distribution
In the section Diversity of CoA and Persuasive Strategy, we further visualize the distribution of persuasive strategy types across different user groups using a stacked bar chart, as shown in Fig.3.
@@ A.4 Prompts of MADS
Prompt Template – User Agent <user_profile> <basic_info> {user_profile} {business_attributes} </basic_info> <personality> Based on the commonly accepted and general {personality_traits} personality traits, please characterize this person’s personality and reflect it in the dialogue. </personality> </user_profile> <task> Based on the above personal profile, please role-play as this person. Note: 1.Make your replies as conversational and colloquial as possible. 2.Only output the dialogue. Do not output any thoughts, inner monologue, or any unnecessary content. </task> Start the task!
Business attributes come from real customer-provided fields, injecting domain-specific prior knowledge. For example, in the insurance marketing scenario:
      {     "health_status": {     "type": "categorical",     "candidates": [     {"value": "Sub-health – mild chronic illness", "p": 0.10},     {"value": "Sub-health – cardiovascular issues", "p": 0.05},     {"value": "Sub-health – digestive system issues", "p": 0.05},     {"value": "Sub-health – endocrine disorders", "p": 0.05},     {"value": "Sub-health – musculoskeletal problems", "p": 0.05},     {"value": "No relevant record", "p": 0.30},     {"value": "Healthy", "p": 0.40}     ]     },     "purchased_insurance_before": {     "type": "boolean",     "candidates": [true, false],     "prior": {"true": 0.5, "false": 0.5}     }    }
@@
A.5 MADS self-play workflow example
[Uncaptioned image][Uncaptioned image][Uncaptioned image]